OpenStudy (anonymous):

If staff salaries were $32,000/month last year and $47,000/month this year, what is the percentage labor cost increase?

OpenStudy (aum):

Percentage increase = Actual increase / original salary * 100 = (47,000 - 32,000) / 32,000 x 100 = ? %

OpenStudy (anonymous):

47%?

OpenStudy (aum):

yeah, 46.9%

OpenStudy (anonymous):

But what if it was $1,500,000 last year and $1,300,000 this year? Would it be: (1,3000,000-1,500,000) /1,3000,000 x 100 = ?

OpenStudy (anonymous):

1,300,000*